OpenStudy (anonymous):

the pilot in an airplane observes the angle of depression of a light directly below his line of sight to be 30.4 degrees. a minute later, its angle of depression is 43.0 degrees. if he is flying horizontally in a straight course at the rate of 150 mph, find the altitude at which he is flying? find his distance from the light at the first point of observation?

OpenStudy (anonymous):

Draw the picture. Fill in the angles you know from the problem description. Draw an altitutde line from the 1st point of observation.

OpenStudy (anonymous):

150 mph = 150/60 = 5/2 = 2.5 miles per minute

OpenStudy (anonymous):

You should have a triangle with angles 30.4,137, and 12.6 and the side with 2.5 miles

OpenStudy (anonymous):

Use the Law of Sines to find "his distance from the light at the 1st point of observation". s/sin(137)= 2.5/sin(12.6) s = 7.82 miles

OpenStudy (anonymous):

Now you can find the altitude using: alt = 7.82*sin(30.4) = 3.96 miles = 20,909 feet

OpenStudy (anonymous):

thank you very much. i'm tired already finding a way to solve the assignment of my daughter.:)